write an equation that satisfies the given conditions. Endpoins of major axis at (9,3) and (-11,3), endpoints of minor axis at (-1,8) and (-1,-2)
Answer by lwsshak3(11628)

write an equation that satisfies the given conditions. Endpoints of major axis at (9,3) and (-11,3), endpoints of minor axis at (-1,8) and (-1,-2)
**
Standard form of equation for an ellipse with horizontal major axis: (x-h)^2/a^2+(y-k)^2/b^2=1, a>b), (h,k) being the (x,y) coordinates of the center:
y-coordinate of center=3
x-coordinate of center=(9+(-11))/2=-2/2=-1 (mid point formula)
center: (-1,3)
length of horizontal major axis=-11 to 9=20=2a
a=10
a^2=100
..
length of minor axis=-2 to 8=10
b=5
b^2=25
..
Equation of given ellipse:
(x+1)^2/100+(y-3)^2/25=1